i don't understand some of the quadratic equation steps, like:
* when to use the two different equations.
*when do i know that there are no solutions
*if i got a fraction for x do i turn it into a decimal?
*if i don't have to turn it into decimal is there a chance that i can get a real big number? how can i work that out?
*when i use the long equation(the one with the square root on it) im looking for the vertex in the parabola or what.. i don't understand?
Please help me im doing kind of bad in Algebra because i don't understand this quadratic equations
Maria Rivera

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Quadratic Equation Form: y = ax^2 + bc + c
---
Quadratic Formula: x = [-b +- sqrt(b^2-4ac)]/(2a)
======================================================
i dont understand some of the quadradic equation steps, like:
* when to use the two different equations.
I think you mean "two different solutions".
Every quadratic has two solutions.
---------
*when do i know that there are no solutions
I think you mean "no Real Number solutions".
If b^2-4ac is negative the equation has no Real Number solutions.
----------------
*if i got a fraction for X do i turn it into a decimanl?
That is your choice. You may change it or you may leave it as a fraction.
--------------------
*if i dont have to turn it into decimal is there a chance that i can get a real big number? how can i work that out?
Yes, solutions may be big, or they may be small.
-----------------------------------------------
*when i use the long equation(the one with the square root on it) im looking for the vertex in the parabola or what.. i dont understand?
You are looking for the zeroes or roots of the equation.
You are looking for the values of "x" which make ax^2+bx+c = 0
You are looking for the x-intercepts of the graph of the equation.
-----------------------------------------------
Cheers,
Stan H.
===================

Answer by richwmiller(17219)


I like your questions
I hope I understand your questions
For every unknown (a different letter), you need another equation.
Let's say you just have x then you only need one equation .
If you have x and y then you need two equations
If you have x, y, and z then you need three equations.
Do you mean when are there no real solutions or no solutions?
There are always solutions to a quadratic equation.
The determinant (b^2-4ac) will tell you what type (real, imaginary, rational, irrational) and how many there are.
Answers can be left in fraction form. They are more exact than the decimal form.
Sometimes the teacher will want an answer in decimal form rounded to one or two places because there are roots with radicals etc.

What do you mean by a really big number? And what is wrong with really big numbers? Wouldn't you like a really big number of dollars?

The vertex will have an x and a y in the point such as (3,4)
the 3 is the x and the 4 is the y
-b/2a is the x value of the vertex when the equation is in the form ax^2+bx+c=0 or ax^2+bx=c
then plug in that value of x to find the y of the point y=ax^2+bx+c or